# Lightweight Learning from Actuation-Space Demonstrations via Flow Matching for Whole-Body Soft Robotic Grasping

### v1

@@ Abstract
Robotic grasping under uncertainty remains a fundamental challenge due to its uncertain and contact-rich nature. Traditional rigid robotic hands, with limited degrees of freedom and compliance, rely on complex model-based and heavy feedback controllers to manage such interactions. Soft robots, by contrast, exhibit embodied mechanical intelligence: their underactuated structures and passive flexibility of their whole body, naturally accommodate uncertain contacts and enable adaptive behaviors. To harness this capability, we propose a lightweight actuation-space learning framework that infers distributional control representations for whole-body soft robotic grasping, directly from deterministic demonstrations using a flow matching model (Rectified Flow), without requiring dense sensing or heavy control loops. Using only 30 demonstrations (less than 8% of the reachable workspace), the learned policy achieves a 97.5% grasp success rate across the whole workspace, generalizes to grasped-object size variations of ±33%, and maintains stable performance when the robot’s dynamic response is directly adjusted by scaling the execution time from 20% to 200%. These results demonstrate that actuation-space learning, by leveraging its passive redundant DOFs and flexibility, converts the body’s mechanics into functional control intelligence and substantially reduces the burden on central controllers for this uncertain-rich task.
@@ I Introduction
Robotic grasping under uncertainty remains a central challenge due to its uncertain and contact-rich nature [1]. These tasks involve high-dimensional, nonlinear, and uncertain dynamics, making them inherently difficult to model and control [2]. Traditional rigid robotic hands, with limited degrees of freedom and low compliance, often exhibit poor adaptability and safety when interacting under such conditions. Soft robots, by contrast, embody mechanical intelligence [3]: their intrinsic redundancy, flexibility, and compliance of their whole body, allow them to passively absorb uncertainty, enabling safe and adaptive physical interaction while reducing reliance on precise centralized control [4]. Despite these morphological advantages, developing motion planning and control strategies for soft robots remains challenging [5]. Their dynamics are high-dimensional and history-dependent, while uncertain contacts resulting from friction [6], material fatigue [7], and geometric constraints [8] further complicate both planning and real-time control.
@@
Existing approaches on the control of soft robots can be broadly categorized into model-based [5] and learning-based methods [9]. Model-based control relies on explicit physical formulations, ranging from piecewise constant curvature (PCC) models [10] and pseudo–rigid-body approximations [8], to geometrically nonlinear beam and rod theories [11, 12], and finite-element or solid-mechanics-based approaches [13, 14]. These controllers offer interpretable descriptions of soft-body mechanics and can often guarantee stability or convergence. However, their practical deployment is hindered by high modeling complexity [11, 12], computationally expensive inverse solutions [13], and large sim-to-real discrepancies arising from material and contact uncertainties [15, 16]. Consequently, purely model-based control remains difficult to scale and lacks robustness when dealing with unmodeled deformations or uncertain interactions.
In contrast, learning-based control circumvents explicit modeling by leveraging data [9]. Artificial neural networks (ANNs) have been widely adopted in soft robotics for tasks such as kinematics, dynamics, and state estimation [17]. For instance, ANNs have been used to approximate the inverse statics of cable-driven soft arms [18], to model forward dynamics for predictive control [19], and to implement imitation learning policies for soft robotic hands trained from human demonstrations [20]. With their universal function approximation capability [21], ANNs can capture nonlinear dynamics directly from data without explicit physical models. However, their accuracy and generalization strongly depend on the amount and quality of training data, and limited demonstrations often constrain their robustness and adaptability [22]. While this suffices for low-level tasks [9], these methods generally fail to cope with robotic manipulation scenarios involving strong dynamics and significant uncertainty [23]. To move beyond supervised imitation, reinforcement learning (RL) has been explored as an alternative that enables autonomous exploration and policy optimization [24]. In soft robotics, RL has been demonstrated in locomotion and basic trajectory tracking, such as soft snake robots [25] and continuum arms [19, 26], yet applications to whole-body grasping remain scarce [17]. Although RL can theoretically surpass expert demonstrations, it is notoriously sample inefficient, requires carefully engineered rewards, and poses safety risks on hardware [7, 24]. More recently, Diffusion Policy (DP)has achieved impressive performance in rigid robot visuomotor control, modeling multimodal action distributions and handling high-dimensional spaces [27]. Extensions include on-device diffusion transformers [28], hierarchical policies for multi-task manipulation [29], and 3D visual representations [30]. Despite their success, diffusion-based methods require multi-step denoising at inference, introducing latency and significant computational cost [27]. Such precise and high-frequency closed-loop corrections may not align with the nature of soft robots, which inherently exhibit nonlinearities, uncertain contacts, and history dependence [3, 6]. In soft robotic systems, embodied mechanical intelligence, arising from redundancy, flexibility, and compliance, already absorbs part of the uncertainty [3, 4], suggesting that increasingly heavy computation may not yield proportional benefits.
@@
These observations motivate a lightweight distributional control framework that generates robust patterns, preserves multimodality, and leverages the embodied mechanical intelligence of soft robots to handle uncertainty. A natural solution is Learning from Demonstrations (LfD), since expert demonstrations inherently encode feasible and multimodal strategies without requiring explicit modeling or reward design [31]. Performing such learning in actuation space may better exploit the robot’s passive flexibility and redundancy, providing a potential pathway for embodied intelligence to emerge through coordinated actuation behaviors. Building on this idea, we propose a lightweight framework for Learning from actuation-space demonstrations that learns control distributions directly aligned with the soft robot’s physical characteristics. Within this framework, we employ the Rectified Flow (RF) model [32], a flow-matching generative approach that efficiently learns control distributions with minimal sampling [33]. We validate the proposed approach on a spiral-shaped soft robot (SpiRob) [8], performing whole-body grasping under uncertainty. The policy, trained with only 30 demonstrations covering less than 8% of the reachable workspace, achieved a 97.5% grasp success rate across the workspace. It also generalized to objects with size variations of $\pm$ 33% and maintained stable grasping performance when the execution duration was scaled between 20% and 200%, demonstrating robust adaptability under uncertainty. The contributions are summarized as follows:
@@ II Problem Statement
II-A Investigated Soft Robot – SpiRob
Spiral robots (SpiRobs) represent a class of bio-inspired continuum manipulators that take advantage of the geometric regularity of logarithmic spirals [8]. Unlike conventional soft arms with uniform structures, SpiRobs exhibit a natural gradient in stiffness and curvature from base to tip, as shown in Fig. 1A. This intrinsic property enables a wide operational workspace and facilitates smooth transitions between different motion primitives such as reaching, wrapping, and grasping. Their structure can be rapidly prototyped through 3D printing of flexible filaments, such as TPU, while actuation is achieved through a minimal number of tendons. This design reduces mechanical complexity and highlights the fundamental challenge of coordinating many coupled degrees of freedom under complex nonlinear body dynamics. These characteristics make SpiRobs an ideal testbed for exploring data-driven approaches in the context of high-dimensional motion generation and control.
II-B Task Objectives
The task suite focuses on object manipulation within a horizontal workspace, designed to exploit the whole-body compliance of SpiRobs. The primary objective is a planar wrapping grasp, where the robot approaches an object and secures it by sequentially uncurling and curling its spiral body.
@@ III-A1 Rectified Flow
Rectified Flow (RF) [32, 33] is a generative flow-matching framework that learns a continuous flow field to transform noise into data, offering a fast and stable alternative to diffusion-based training [34]. The overall training and inference pipeline of the RF model is illustrated in Fig. 2. Originally proposed in computer vision, RF has demonstrated a strong ability to capture complex and multimodal distributions, enabling more robust and generalizable control than conventional imitation or reinforcement learning [35]. Let $U_{1}\sim\Pi_{\text{data}}$ denote an expert training sample, and $U_{0}\sim\mathcal{N}(0,\mathbb{I})$ a Gaussian noise sample. RF learns a vector field $v_{\tau}(U_{\tau})$ that continuously transports $U_{0}$ to $U_{1}$ over the normalized flow time $\tau\in[0,1]$, governed by the ordinary differential equation:
@@ III-B1 Learning from Actuation-Space Demonstration
Soft Robot Dynamics. Soft robots, being underactuated systems, particularly continuum robots, possess a continuous backbone with a large number degrees of freedom but only a limited number of actuators [10]. The studied SpiRob (Fig. 1A) is modeled in MuJoCo using the concept of pseudo-rigid-body model [36], where the continuous structure is discretized into a chain of rigid segments connected by revolute joints. The resulting dynamics are formulated as
@@
where $q$, $\dot{q}$, $\ddot{q}\in\mathbb{R}^{n}$ denote the generalized displacement, velocity, and acceleration, respectively. $M(q)$ represents the inertia matrix, $C(q,\dot{q})$ the Coriolis and centrifugal effects, $G(q)$ the gravity vector, $K(q)$ the stiffness term capturing backbone flexibility, and $D(q,\dot{q})$ the damping term reflecting internal frictional effects. The actuation mapping matrix $H_{c}(q)\in\mathbb{R}^{m\times n}$ is generally rectangular with $m<n$, indicating underactuation and a non-bijective mapping between actuation and configuration. According to the definition in Section III-A2, $U^{*}$ explicitly denotes the system input as a continuous function in (6). The investigated SpiRob has a configuration dimension of $n=24$ with $m=2$ active tendons, clearly exhibiting underactuation. Tendon–robot friction contributes to multiple stable postural configurations [6], while external contact and internal self contact, as shown in Fig. 1A, represented by $H_{e}^{\mathrm{T}}(q)u_{e}$, introduce the strong nonlinear behaviors  [8]. Each tendon’s geometric displacement is then defined as
$\small\Delta l_{i}=l_{i}(q)-l_{i}(q_{0}),\ i=1,2.$ (7)
where $l_{i}(q)$ denotes the tendon length at configuration $q$, and $l_{i}(q_{0})$ its reference length at the undeformed configuration $q_{0}=\mathbf{0}$.
@@
Rigid Robot Dynamics. In contrast, rigid robots are fully actuated systems, where the generalized coordinates $q(t)$ and the actuation input $U^{*}(t)$ are uniquely related through:
$\small M(q)\ddot{q}+C(q,\dot{q})+G(q)+D(q,\dot{q})=H(q)_{c}^{\mathrm{T}}U^{*}+H_{e}^{\mathrm{T}}(q)u_{e}$ (8)
@@
where $H(q)_{c}\in\mathbb{R}^{n\times n}$ is a square full-rank matrix. This bijective mapping between actuation and configuration spaces allows demonstrations to be directly performed in the configuration space. During demonstration, dragging the end-effector along a desired trajectory $q_{e}(t)$ uniquely determines the joint configurations $q(t)$. The resulting $q(t)$ is directly converted by the low-level joint displacement controller into the control signal $U^{*}(t)$, ensuring deterministic and repeatable actuation.
Comparison and Motivation. Rigid robots, benefiting from full actuation, allow precise configuration-space teaching and deterministic control. They are inherently suited for high-precision and repeatable tasks in well-defined environments but become limited when facing dynamic interactions with high uncertainty, such as grasping, unless equipped with high-frequency, model-based feedback control. In contrast, soft robots are underactuated and possess high degrees of freedom. Their configuration $q$ cannot be directly controlled due to underactuation, yet this very property, together with passive compliance, allows the actuation space to fully express their embodied mechanical intelligence. Learning in the actuation space therefore exploit and amplify these intrinsic features, enabling adaptive behaviors to emerge through coordinated actuation. Consequently, demonstrations for soft robots are conducted in the actuation space, where each sample is represented as a control–task pair. Therefore, demonstrations for soft robots are instead performed in the actuation space, where each sample is represented as a control–task pair:
$\small(U^{*}(t),\,o)$ (9)
with $U^{*}(t)$ denoting the demonstrated control sequence and $o$ representing the task condition. Training in this actuation space allows soft robots to directly learn the correspondence between actuation inputs and task outcomes, avoiding the need for explicit and accurate inverse modeling from configuration to actuation and enabling robust adaptation to contact-rich environments (Fig. 1B). A concise comparison of rigid and soft robots in terms of task, configuration, and actuation spaces is summarized in Table I.
TABLE I:
Robot Task Space Configuration Space Actuation Space
Rigid Precise Deterministic Deterministic
Soft Uncertain Adaptive Distributional
Algorithm 1 Rectified Flow for Learning Actions of SpiRob
1def train_step(batch):
2 state, target_action = batch
3 cond = state_encoder(state)
4 tau = rand() # normalized flow time
5 U0 = randn() # Gaussian noise
6 U1 = target_action # expert control
7 target_flow = U1 - U0
8 U_tau = U0 + tau * target_flow
9 pred_flow = model(U_tau, tau, cond)
10 loss = mse(pred_flow, target_flow)
11 return loss
12
13def inference_step(state):
14 cond = state_encoder(state)
15 U = randn()
16 for k in range(k_steps):
17 tau = k / k_steps
18 d_tau = 1 / k_steps
19 pred_flow = model(U, tau, cond)
20 U = U + pred_flow * d_tau
21 return U
@@ III-B2 Model Architecture
To learn the control strategy of SpiRobs in whole-body grasping under uncertainty via RF, we develop a MuJoCo-based training framework forming the foundation of our model architecture. As illustrated in Fig. 2A, the proposed rectified-flow model, conditioned on the task condition which is object state $o\in\mathbb{R}^{C}$ in this studed case, learns an approximated flow field $v_{\theta}(U_{\tau},\tau,o)$ that transforms Gaussian noise $U_{0}\in\mathbb{R}^{N\times 2}$ into the expert control sequence $U_{1}\in\mathbb{R}^{N\times 2}$ for the studied SpiRob, where $C$ denotes the dimension of the object state and $N$ the number of timesteps. The detailed implementation procedure is summarized in Algorithm 1.
@@
For each normalized flow time $\tau\in[0,1]$, the model generates scale, shift, and gate factors to modulate intermediate feature representations within transformer blocks:
$\small h^{\prime}=\gamma\odot(\alpha\odot h+\beta)$ (10)
where $\alpha$, $\beta$, and $\gamma$ denote the scale, shift, and gate factors, respectively. This modulation mechanism adaptively regulates feature magnitudes and ensures consistent evolution of the internal flow representation along $\tau$. An Action Decoder, implemented as a convolutional neural network (CNN), maps the latent features back to the target control sequence space.

### v2

@@ Abstract
Robotic grasping under uncertainty remains a fundamental challenge due to its uncertain and contact-rich nature. Traditional rigid robotic hands, with limited degrees of freedom and compliance, rely on complex model-based and heavy feedback controllers to manage such interactions. Soft robots, by contrast, exhibit embodied mechanical intelligence: their underactuated structures and passive flexibility of their whole body naturally accommodate uncertain contacts and enable adaptive behaviors. To harness this capability, we propose a lightweight actuation-space learning framework that infers distributional control representations for whole-body soft robotic grasping directly from deterministic demonstrations using a flow matching model (Rectified Flow), without requiring dense sensing or heavy control loops. Trained with only 30 demonstrations covering less than $8\%$ of the reachable workspace, the learned policy achieved a $97.5\%$ grasp success rate over 1000 trials in simulation. In real-world experiments on 50 uniformly distributed targets, the policy achieved a $100\%$ success rate, generalized to object size variations from $33\%$ to $100\%$, and remained stable under execution-time scaling from $20\%$ to $200\%$. These results demonstrate that actuation-space learning effectively embeds mechanical intelligence into control, significantly reducing reliance on centralized computation for grasping under uncertainty.
@@ I Introduction
Robotic grasping in unstructured environments is inherently contact-rich and uncertain, making it difficult to model and control reliably [1]. Rigid hands, with low compliance, often exhibit limited adaptability and safety under such uncertainty. Soft robots instead exploit whole-body compliance and redundancy to passively absorb disturbances and tolerate contact variations [2], reducing reliance on precise centralized control, often described as embodied/mechanical intelligence [3]. A large body of soft grasping and manipulation research has focused on tendon-driven soft grippers/hands and cable-driven continuum manipulators, which enable adaptive interaction through compliance and underactuated actuation [4]. Yet planning and control remain challenging because soft-body dynamics are high-dimensional and history-dependent, and friction, material fatigue, and geometric constraints can dominate behavior during interaction [5].
@@
Existing approaches on the control of soft robots can be broadly categorized into model-based [6] and learning-based methods [7]. Model-based control relies on explicit physical formulations, ranging from piecewise constant curvature (PCC) models [8] and pseudo–rigid-body approximations [5], to geometrically nonlinear beam and rod theories [9], and finite-element or solid-mechanics-based approaches [10]. These controllers offer interpretable descriptions of soft-body mechanics and can often guarantee stability or convergence. However, their practical deployment is hindered by high modeling complexity [9], computationally expensive inverse solutions [10], and large sim-to-real discrepancies arising from material and contact uncertainties [11]. Consequently, purely model-based control remains difficult to scale and lacks robustness when dealing with unmodeled deformations or uncertain interactions.
@@
In contrast, learning-based control circumvents explicit modeling by leveraging data [7]. Artificial neural networks (ANNs) have been widely adopted in soft robotics for tasks such as kinematics, dynamics, and state estimation [12]. For instance, ANNs have been used to approximate the inverse statics of cable-driven soft arms [13], to model forward dynamics for predictive control [14], and to implement imitation learning policies for soft robotic hands trained from human demonstrations [15]. With their universal function approximation capability [16], ANNs can capture nonlinear dynamics directly from data without explicit physical models. However, their accuracy and generalization strongly depend on the amount and quality of training data, and limited demonstrations often constrain their robustness and adaptability [17]. While this suffices for low-level tasks [7], these methods generally fail to cope with robotic manipulation scenarios involving strong dynamics and significant uncertainty [18]. To move beyond supervised imitation, reinforcement learning (RL) has been explored as an alternative that enables autonomous exploration and policy optimization [19]. In soft robotics, reinforcement learning (RL) has been mainly demonstrated in locomotion and basic trajectory tracking, such as soft snake robots [20] and continuum arms [14, 21]. Recent works have begun to explore interaction-rich manipulation with soft robotic arms, including pushing behaviors [22] and RL-based characterization of continuum arms [23]. Nevertheless, applications of RL to whole-body grasping with strong generalization under uncertainty remain scarce [12].
Although RL can theoretically surpass expert demonstrations, it is notoriously sample inefficient, requires carefully engineered rewards, and poses safety risks on hardware [19]. Learning from demonstrations (LfD) reduces the reliance on costly exploration [24]. However, for highly redundant soft robots, the key challenge shifts to how generalization can be achieved from a limited set of demonstrations [7]. More recently, Diffusion Policy (DP) has achieved impressive performance in rigid robot visuomotor control by modeling multimodal action distributions and handling high-dimensional spaces [25]. Extensions include on-device diffusion transformers [26], hierarchical policies for multi-task manipulation [27], and 3D visual representations [28], further enhancing their generalization across tasks and environments. Despite their success, diffusion-based methods require multi-step denoising at inference, introducing latency and significant computational cost [25]. Such precise, high-frequency closed-loop corrections may not align with the embodied nature of soft robots, where redundancy and compliance inherently mitigate uncertainty [3], making heavy per-timestep state observation and diffusion-based inference computationally inefficient.
These observations motivate a lightweight, diffusion-inspired flow-matching approach that generalizes from limited demonstrations while avoiding heavy per-timestep inference. Specifically, we employ the Rectified Flow (RF) model [29], a flow-matching generative framework that efficiently learns distributional control policies from minimal demonstrations under task conditions [30]. We validate the proposed approach on a spiral-shaped soft robot (SpiRob) [5] for whole-body grasping under uncertainty. Trained with only 30 demonstrations covering less than $8\%$ of the reachable workspace, the learned policy achieves a $97.5\%$ grasp success rate over 1000 simulation trials. In real-world experiments, 50 uniformly distributed targets demonstrate a $100\%$ grasp success rate, confirming robust adaptability under uncertainty. The approach further generalizes to object size variations ranging from $33\%$ to $100\%$, extends to diverse object shapes, and remains stable when the execution duration is scaled between $20\%$ and $200\%$. The main contributions of this work are summarized as follows:
@@ II Problem Statement
II-A Task Objectives
We study whole-body grasping in a horizontal plane using a soft continuum arm. Given a cylindrical object at a specified planar position, the robot must approach the target and achieve enclosure by uncurling and then curling its body. For repeatable evaluation and to isolate grasp synthesis from uncontrolled object sliding/rolling, the object is constrained (fixed) during the wrapping phase; this also reflects supported or handover-like settings where the object is externally stabilized.
II-B Investigated Soft Robot – SpiRob
Spiral robots (SpiRobs) are a class of bio-inspired continuum manipulators that exploit the geometric regularity of logarithmic spirals [5]. Unlike many tendon-/cable-driven continuum robots that obtain dexterity through multi-tendon and/or multi-segment actuation [4], SpiRobs achieve a wide workspace and smooth transitions among motion primitives such as reaching, wrapping, and grasping with only two tendons on a logarithmic-spiral body (Fig. 1A). This minimalist design reduces structural complexity and facilitates rapid fabrication (e.g., 3D-printed TPU), making SpiRobs an attractive platform for whole-body grasping. However, the same minimal actuation also introduces substantial control challenges. Due to strong underactuation and distributed contact, friction and hysteresis can significantly alter global deformation and interaction forces, making the actuation-to-interaction mapping highly nonlinear and history-dependent [31]. This underactuated, contact-rich complexity motivates learning coordinated actuation patterns directly from demonstrations.
@@ III-A1 Rectified Flow
Rectified Flow (RF) [29, 30] is a generative flow-matching framework that learns a continuous flow field to transform noise into data, offering a fast and stable alternative to diffusion-based training [32]. The overall training and inference pipeline of the RF model is illustrated in Fig. 2. Let $U_{1}\sim\Pi_{\text{data}}$ denote an expert training sample, and $U_{0}\sim\mathcal{N}(0,\mathbb{I})$ a Gaussian noise sample. RF learns a vector field $v_{\tau}(U_{\tau})$ that continuously transports $U_{0}$ to $U_{1}$ over the normalized flow time $\tau\in[0,1]$, governed by the ordinary differential equation:
@@ III-B1 Learning from Actuation-Space Demonstration
Soft Robot Dynamics. Soft robots are inherently underactuated systems, characterized by a continuous backbone with a high-dimensional configuration space but only a limited number of actuators [8]. The studied SpiRob (Fig. 1A) is modeled in MuJoCo using a pseudo-rigid-body formulation [33], where the continuum structure is discretized into a chain of rigid segments connected by revolute joints. Its dynamics are expressed as
@@
where $q$, $\dot{q}$, and $\ddot{q}\in\mathbb{R}^{n}$ denote the generalized displacement, velocity, and acceleration, respectively. Here, $M(q)$ denotes the inertia matrix, $C(q,\dot{q})$ captures Coriolis and centrifugal effects, $G(q)$ represents gravity, $K(q)$ models the elastic stiffness of the backbone, and $D(q,\dot{q})$ aggregates internal damping and frictional effects. External and self-contact forces are represented by $H_{e}^{\mathrm{T}}(q)u_{e}$. Crucially, the actuation mapping matrix $H_{c}(q)\in\mathbb{R}^{m\times n}$ is rectangular with $m<n$, reflecting strong underactuation. For the investigated SpiRob, only $m=2$ tendons are available to actuate a system with $n=24$ degrees of freedom. Consequently, the set of configurations reachable through actuation constitutes only a subset of the configuration space [34]. As a result, a directly specified or demonstrated configuration trajectory $q(t)$ may admit no corresponding actuation input $U^{*}(t)$ capable of realizing it. Therefore, for underactuated soft robots, configuration trajectories $q(t)$ are not valid control variables for demonstration or inversion, but instead emerge as outcomes of actuation, system dynamics, and interactions with the environment.
@@
Rigid Robot Dynamics. In contrast, rigid robots are typically fully actuated systems, where the generalized coordinates $q(t)$ and the actuation input $U^{*}(t)$ are well defined through joint-space feedback on $q$:
$\small M(q)\ddot{q}+C(q,\dot{q})+G(q)+D(q,\dot{q})=H_{c}^{\mathrm{T}}(q)U^{*}+H_{e}^{\mathrm{T}}(q)u_{e},$ (7)
@@
with $H_{c}(q)\in\mathbb{R}^{n\times n}$ being a square, full-rank matrix. This full-rank actuation mapping allows demonstrations to be naturally specified in the configuration space. During demonstration, dragging the end-effector along a desired trajectory $q_{e}(t)$ uniquely determines the joint configuration $q(t)$, which is tracked by a joint-space displacement feedback controller to generate the corresponding actuation input $U^{*}(t)$, resulting in deterministic and repeatable control (Fig. 1B).
Comparison and Motivation. As discussed above, configuration-space demonstrations, which are natural and effective for fully actuated rigid robots, do not directly transfer to soft robots. Due to inherent underactuation and redundancy, soft robots cannot be reliably controlled or demonstrated through configuration trajectories. Instead, the actuation space provides a more appropriate interface for demonstration, as it directly captures the robot’s embodied mechanical intelligence arising from compliance and redundancy. Learning from demonstrations in the actuation space therefore leverages these intrinsic properties, enabling adaptive behaviors without requiring explicit inverse mappings from configuration to actuation. Accordingly, demonstrations for soft robots are represented as actuation–task pairs
$\small(U^{*}(t),\,o),$ (8)
where $U^{*}(t)$ denotes the demonstrated control sequence and $o$ represents the task condition (Cartesian coordinates of the object to be grasped). Training in the actuation space allows the robot to learn the correspondence between actuation inputs and task outcomes, supporting robust adaptation in contact-rich and uncertain environments (Fig. 1B).
@@ III-B2 Model Architecture
To learn the control strategy of SpiRobs in whole-body grasping under uncertainty via RF, we develop a MuJoCo-based training framework forming the foundation of our model architecture. As illustrated in Fig. 2A, the proposed rectified-flow model, conditioned on the task condition (the object state $o\in\mathbb{R}^{2}$ in this studied case), learns an approximated flow field $v_{\theta}(U_{\tau},\tau,o)$ that transforms Gaussian noise $U_{0}\in\mathbb{R}^{N\times 2}$ into the expert control sequence $U_{1}\in\mathbb{R}^{N\times 2}$ for the studied SpiRob, where $N$ is the number of timesteps. During execution, the trained policy performs a single open-loop inference at the beginning of the grasp, taking $o$ as input and generating the the expert control sequence $U_{1}$ without online replanning.
@@
For each normalized flow time $\tau\in[0,1]$, the model generates scale, shift, and gate factors to modulate intermediate feature representations within transformer blocks: $\small h^{\prime}=\gamma\odot(\alpha\odot h+\beta)$ where $\alpha$, $\beta$, and $\gamma$ denote the scale, shift, and gate factors, respectively. This modulation mechanism adaptively regulates feature magnitudes and ensures consistent evolution of the internal flow representation along $\tau$. An Action Decoder, implemented as a convolutional neural network (CNN), maps the latent features back to the target control sequence space.